# Can Transformers Learn Full Bayesian Inference in Context?

### v1

@@ Abstract
Transformers have emerged as the dominant architecture in the field of deep learning, with a broad range of applications and remarkable in-context learning (ICL) capabilities. While not yet fully understood, ICL has already proved to be an intriguing phenomenon, allowing transformers to learn in context—without requiring further training. In this paper, we further advance the understanding of ICL by demonstrating that transformers can perform full Bayesian inference for commonly used statistical models in context. More specifically, we introduce a general framework that builds on ideas from prior fitted networks and continuous normalizing flows which enables us to infer complex posterior distributions for methods such as generalized linear models and latent factor models. Extensive experiments on real-world datasets demonstrate that our ICL approach yields posterior samples that are similar in quality to state-of-the-art MCMC or variational inference methods not operating in context.
00footnotetext: Correspondence to Arik Reuter: arik.reuter@campus.lmu.de
@@ 1 Introduction
Furthermore, ICL has recently shown remarkable promise for regression and classification tasks involving tabular data, with tabular prior-data fitted networks (TabPFNs) dominating regression and classification benchmarks alongside minimal prediction time (Hollmann et al., 2022, 2025). While the internet serves as a suitable source for the massive data needed to train in-context learners on text, TabPFNs demonstrate that training on purely synthetic data facilitates the development of in-context learners for tabular data.
@@
Figure 1: (a) An LLM generates a summary $s_{1},s_{2},\ldots$ of a text $t_{1},t_{2},\ldots,t_{K}$ through autoregressive sampling while referring to the context using masked self-attention. (b) A dataset ${\bm{x}}$ is processed with a transformer encoder. Subsequently cross attention allows to generate samples from the posterior conditioned on ${\bm{x}}$ in context using a diffusion transformer (decoder) and flow matching.
@@
While PFNs perform Bayesian inference, they target a univariate, typically discrete, posterior predictive distribution. In numerous applications, however, high-dimensional and continuous posteriors $P^{{\bm{z}}|{\bm{x}}}$ of (latent) variables ${\bm{z}}$ given data ${\bm{x}}$ play a key role [^1] We do not assume any specific form of ${\bm{z}}$. That is, there can be a single ${\bm{z}}_{j}$ associated with each data point ${\bm{x}}_{j}$ in ${\bm{x}}$, but the case where a single “global” ${\bm{z}}$ governs the behavior of each ${\bm{x}}_{j}$ in ${\bm{x}}$ is equally included in this notation.. This includes areas such as healthcare (Kyrimi et al., 2021; Abdullah et al., 2022; Etzioni & Kadane, 1995), physics (Gebhard et al., 2025; Brehmer & Cranmer, 2022; Dax et al., 2024), and neuroscience (Lueckmann et al., 2017; Sohn & Narain, 2021). We use the notion of full Bayesian inference for methods yielding potentially complex and high-dimensional posterior distributions—in contrast to, for instance, methods that yield only the posterior predictive or point estimates of the posterior as, for example Hollmann et al. (2022). However, performing full Bayesian inference can be challenging, even for relatively simple models such as generalized linear models (GLMs; Nelder & Wedderburn, 1972). Two common issues when performing full Bayesian inference include (a) slow inference time, particularly when using sampling-based methods (Sommer et al., 2025, 2024), and (b) model misspecification. Although potentially restrictive modeling assumptions are often necessary to make Bayesian inference efficient or even feasible, they can lead to suboptimal predictive performance (Wang & Blei, 2019; Walker, 2013).
@@
In this paper, we address the following question: Can we leverage in-context learning to effectively perform full Bayesian inference? In doing so, we aim to obtain an in-context learner that can perform the mapping ${\bm{x}}\mapsto P^{{\bm{z}}|{\bm{x}}}$, and, analogous to LLMs, (a) allows for the rapid generation of samples from a posterior of interest during deployment and (b) can flexibly adapt to a broad range of inputs, thereby overcoming issues arising from model misspecification.
The remainder of this paper is structured as follows: Section 2 discusses how ideas from the in-context learning literature, amortized inference and simulation-based inference relate to each other and form the foundation for performing full Bayesian inference in context. Section 3 explains how the generative nature of many probabilistic models allows sampling synthetic data we can use to train the in-context learner via normalizing flows. Then, we present the results of our in-context learning approach on extensive real-world and synthetic datasets in Section 4. Finally, we discuss the challenges and the transformative potential of in-context learning for full Bayesian inference in Section 5.
@@
    We then analyze the efficacy of our approach for GLMs and latent factor models. For these applications, we show that including the “prior” used for TabPFNs results in reliably inferring posterior distributions on real-world data.
@@
    In a variety of experiments, we demonstrate that this approach yields posterior samples that are very similar to those from a Hamiltonian Monte Carlo sampler. Furthermore, we find that the quality of the samples, when compared to various popular VI techniques that do not operate in context, is preferable. [^2] The source code for this paper is available at https://github.com/ArikReuter/ICL_for_Full_Bayesian_Inference.
@@ 2 Related Work
Beyond the perspective of prior-data fitted networks, the contribution of this work can be summarized from the viewpoints of recent work on in-context learning, amortized Bayesian inference, and, in particular, simulation-based inference.
@@
ICL is a special case of meta-learning (Hospedales et al., 2021) characterized by using a large pre-trained model in order to learn from a context dataset without explicitly updating task-specific parameters. Several recent lines of work investigate the in-context learning capabilities of transformers (Garg et al., 2022; Ahuja et al., 2023; Wang et al., 2024; Chan et al., 2022). For instance, Garg et al. (2022) show that a model similar to GPT-2 can implicitly implement various interesting function classes in context. More specifically, the model learns to reproduce the predictions of different statistical models such as (sparse) linear functions, decision trees, and even two-layer neural networks. This approach can be extended to multiple families of functions and even mixtures of tasks (Ahuja et al., 2023). However, the results by Garg et al. (2022) and Ahuja et al. (2023) are restricted to relatively small problem scales and scalar-valued predictions instead of multivariate posterior distributions. Additionally, the experiments are conducted exclusively on simulated data. In contrast, our results show that (large) transformer models can effectively learn multivariate posterior distributions in context on real-world datasets.
@@
Amortized inference is a central paradigm in the field of variational inference (Kingma, 2013; Zhai et al., 2018; Kim et al., 2018; Margossian & Blei, 2023). The central idea here is to model the posterior distribution $P^{{\bm{z}}|{\bm{x}}}$ of latent variables ${\bm{z}}$ given a dataset ${\bm{x}}$ via a factorized density $p({\bm{z}}|{\bm{x}})\approx\prod_{j=1}^{K}q_{\theta}({\bm{z}}_{j}|h_{\bm{\phi}}({\bm{x}}_{j}))$. In contrast to our more general assumption, each datapoint ${\bm{x}}_{j}$ in ${\bm{x}}$ is assumed to have a corresponding latent variable ${\bm{z}}_{j}$. While the parameter $\theta$ determines global aspects of the variational distribution, the function $h_{\bm{\phi}}$ is shared for all ${\bm{x}}_{j}$ and thus amortized across a dataset ${\bm{x}}$. For example, variational autoencoders (Kingma, 2013; Rezende et al., 2014) and neural processes (Garnelo et al., 2018a, b; Rudner et al., 2018) are important model classes based on amortized inference.
In comparison, our ICL approach amortizes its parameters on the level of datasets, such that a single functional relationship is learned for a set $\mathcal{D}\subset(\mathcal{X}\times\mathcal{Z})^{N}$ of datasets. From this point of view, $\mathcal{D}=\left\{\left({\bm{x}}^{(i)},{\bm{z}}^{(i)}\right)\right\}_{i=1}^{N}$ comprising $N$ datasets ${\bm{x}}^{(i)}\in\mathcal{X}$ and the corresponding latent variables ${\bm{z}}^{(i)}\in\mathcal{Z}$ can be seen as a “meta-dataset” for which we perform amortized inference. Furthermore, unlike amortized variational inference, we do not use the notion of an evidence lower bound (Blei et al., 2017) or even the Kullback-Leibler divergence to learn the posterior distribution, but rather utilize ideas that also appear in the context of simulation-based inference.
@@
Analogously to latent variable models, some scientific simulations, for instance in neuroscience or astrophysics (Fan & Markram, 2019; Schmit & Pritchard, 2018), allow to draw samples from the joint distribution $P^{{\bm{x}},{\bm{z}}}$ of data and latent variable of interest. Amortized posterior inference in this context is referred to as simulation-based inference (SBI; Cranmer et al., 2020). Several recent approaches focus on using neural networks to directly infer aspects of the likelihood $p({\bm{x}}|{\bm{z}})$, the posterior $P^{{\bm{z}}|{\bm{x}}}$ or the joint distribution $P^{{\bm{x}},{\bm{z}}}$ in the aforementioned simulation cases. More specifically, techniques based on discrete normalizing flows (Dax et al., 2021) or flow-matching (Wildberger et al., 2024) are used to approximate the posterior $P^{{\bm{z}}|{\bm{x}}}$, while Gloeckler et al. (2024) propose to use a transformer-based diffusion model in order to approximate the joint distribution $P^{{\bm{x}},{\bm{z}}}$.
From a simulation-based inference viewpoint, we demonstrate that sample-based posterior estimation (Dax et al., 2021) can be used for full Bayesian inference in complex scenarios arising in commonly used latent variable models, and demonstrate the effectiveness of this approach on real-world datasets.
@@ 3 In-context learning for full Bayesian inference
We propose a simple and effective solution based on ideas from ICL, which can be seen as conducting amortized inference on a dataset level. Training a model on a potentially unlimited amount of synthetic datasets yields an in-context learner that can not only approximate a vast, almost arbitrarily large, class of distributions but is also highly efficient when used for sampling. Furthermore, this does not incur any issues with overly or insufficiently flexible distribution assumptions as in VI. In the following, we describe a sufficient general condition, as well as a specific framework that allows to train probabilistic in-context learners on simulated data.
@@
The idea underlying the proposed approach is founded on two observations relating to full Bayesian inference and the working principle of PFNs: First, many Bayesian models have a generative formulation that allows the simulation of arbitrarily large amounts of training samples from the joint distribution $P^{{\bm{x}},{\bm{z}}}$. We assume that samples from $P^{{\bm{x}},{\bm{z}}}$ comprise a dataset $\smash{{\bm{x}}=\left\{{\bm{x}}_{j}\right\}_{j=1}^{K}}$ containing $K$ samples ${\bm{x}}_{j}\in\mathcal{X}$ and a corresponding (latent) variable ${\bm{z}}\in\mathcal{Z}$ [^3] We do not assume any specific form of ${\bm{z}}$. That is, there can be a single ${\bm{z}}_{j}$ associated with each data point ${\bm{x}}_{j}$ in ${\bm{x}}$, but the case where a single “global” ${\bm{z}}$ governs the behavior of each ${\bm{x}}_{j}$ in ${\bm{x}}$ is equally included in this notation. This joint distribution $P^{{\bm{x}},{\bm{z}}}$ corresponds to the “prior” in PFNs and allows the training of a large neural network that implicitly learns to perform Bayesian inference. Second, Bayesian inference is especially useful for smaller datasets ${\bm{x}}$ that can be processed in a single forward pass. This makes an entire dataset a viable context for Bayesian ICL.
@@
More specifically, the central goal is to develop a method allowing to infer the posterior distribution $P^{{\bm{z}}|{\bm{x}}}$ of latent variables ${\bm{z}}\in\mathcal{Z}$, given observations ${\bm{x}}\in\mathcal{X}$ using ICL. From a supervised-learning perspective, we thus aim to directly learn the mapping $f_{0}:\mathcal{X}\rightarrow\mathcal{M}(\mathcal{Z}),{\bm{x}}\mapsto P^{{\bm{z}}|{\bm{x}}}$, where $\mathcal{M}(\mathcal{Z})$ is the space of all probability measures. Therefore, we want a model $f_{\theta}({\bm{x}})=Q^{{\bm{z}}|{\bm{x}}}_{\theta}$ for the posterior to be as close as possible to the true posterior $P^{{\bm{z}}|{\bm{x}}}=f_{0}({\bm{x}})$ We measure “closeness” w.r.t. some divergence $d:\mathcal{M}(\mathcal{Z})\times\mathcal{M}(\mathcal{Z})\to[0,\infty)$. When considering the expected divergence over data samples ${\bm{x}}\sim P^{{\bm{x}}}$, this gives rise to the following objective:
$\mathcal{R}_{\theta}\vcentcolon=\mathbb{E}_{{\bm{x}}\sim p({\bm{x}})}\left[d\left(f_{\theta}({\bm{x}}),f_{0}({\bm{x}})\right)\right]=\mathbb{E}_{{\bm{x}}\sim p({\bm{x}})}\left[d\left(Q^{{\bm{z}}|{\bm{x}}}_{\theta},P^{{\bm{z}}|{\bm{x}}}\right)\right].$ (1)
@@
Note that we use the notion of a divergence $d$ loosely to refer to any measure of similarity of two distributions. Although $\mathcal{R}_{\theta}$ itself is usually intractable, specific choices of $d$ and the use of the joint distribution $P^{{\bm{x}},{\bm{z}}}$ make Eq. 1 accessible via

### v2

@@ Abstract
Transformers have emerged as the dominant architecture in the field of deep learning, with a broad range of applications and remarkable in-context learning (ICL) capabilities. While not yet fully understood, ICL has already proved to be an intriguing phenomenon, allowing transformers to learn in context—without requiring further training. In this paper, we further advance the understanding of ICL by demonstrating that transformers can perform full Bayesian inference for commonly used statistical models in context. More specifically, we introduce a general framework that builds on ideas from prior fitted networks and continuous normalizing flows and enables us to infer complex posterior distributions for models such as generalized linear models and latent factor models. Extensive experiments on real-world datasets demonstrate that our ICL approach yields posterior samples that are similar in quality to state-of-the-art MCMC or variational inference methods that do not operate in context. The source code for this paper is available at https://github.com/ArikReuter/ICL_for_Full_Bayesian_Inference
In-Context Learning, Prior-Data Fitted Networks, Bayesian Inference
@@ 1 Introduction
Furthermore, ICL has recently shown remarkable promise for regression and classification tasks involving tabular data, with tabular prior-data fitted networks (TabPFNs) dominating benchmarks alongside minimal prediction time (Hollmann et al., 2022, 2025; Hoo et al., 2024; Robertson et al., 2024). While the internet serves as a suitable source for the massive data needed to train in-context learners on text, TabPFNs demonstrate that training on purely synthetic data facilitates the development of in-context learners for tabular data.
@@
Figure 1: (a) An LLM generates a summary $s_{1},s_{2},\ldots$ of a text $t_{1},t_{2},\ldots,t_{K}$ through autoregressive sampling while referring to the context using masked self-attention. (b) A dataset $\bm{x}$ is processed with a transformer encoder. Subsequently, cross attention allows generating samples from the posterior conditioned on $\bm{x}$ in context using a diffusion transformer (decoder). The samples are generated by solving a neural differential equation defining a continuous normalizing flow.
@@
While PFNs perform Bayesian inference, they target a univariate, typically discrete, posterior predictive distribution. In numerous applications, however, high-dimensional and continuous posteriors $P^{\bm{z}|\bm{x}}$ of (latent) variables $\bm{z}$ given data $\bm{x}$ play a key role. [^1] We do not assume any specific form of $\bm{z}$. That is, there can be a single $\bm{z}_{j}$ associated with each data point $\bm{x}_{j}$ in $\bm{x}$, but the case where a single “global” $\bm{z}$ governs the behavior of each $\bm{x}_{j}$ in $\bm{x}$ is equally included in this notation. This includes areas such as healthcare (Kyrimi et al., 2021; Abdullah et al., 2022; Etzioni & Kadane, 1995), physics (Gebhard et al., 2025; Brehmer & Cranmer, 2022; Dax et al., 2024), and neuroscience (Lueckmann et al., 2017; Sohn & Narain, 2021). We use the notion of full Bayesian inference for methods yielding potentially complex and high-dimensional posterior distributions—in contrast to, for instance, methods that yield only the posterior predictive or point estimates of the posterior as, for example Hollmann et al. (2022). However, performing full Bayesian inference can be challenging, even for relatively simple models such as generalized linear models (GLMs; Nelder & Wedderburn, 1972). Two common issues when performing full Bayesian inference include (a) slow inference time, particularly when using sampling-based methods (Sommer et al., 2025, 2024), and (b) model misspecification. Although potentially restrictive modeling assumptions are often necessary to make Bayesian inference efficient or even feasible, they can lead to suboptimal predictive performance (Wang & Blei, 2019; Walker, 2013).
@@
In this paper, we address the following question: Can we leverage in-context learning to effectively perform full Bayesian inference? In doing so, we aim to obtain an in-context learner that can perform the mapping $\bm{x}\mapsto P^{\bm{z}|\bm{x}}$ for a specific probabilistic model, and, analogous to LLMs, (a) allows for the rapid generation of samples from a posterior of interest during deployment and (b) can flexibly adapt to a broad range of inputs, thereby overcoming issues arising from model misspecification. More specifically, our approach combines a TabPFN encoder (Hollmann et al., 2022) and a diffusion transformer-decoder (Peebles & Xie, 2023) that is trained via flow matching (Lipman et al., 2022).
We present the results of our in-context learning approach on extensive real-world and synthetic datasets in Section 4 and discuss the challenges and the transformative potential of in-context learning for full Bayesian inference in Section 5.
@@
    We then analyze the efficacy of our approach for GLMs and latent factor models, namely Gaussian mixture models (GMMs) and factor analysis (FA). For these applications, we show that including the “prior” used for TabPFNs results in reliably inferring posterior distributions on real-world data.
@@
    In a variety of experiments, we demonstrate that this approach yields posterior samples that are very similar to those from a Hamiltonian Monte Carlo sampler. Furthermore, we find that the quality of the samples from our ICL approach is preferable, when compared to various popular VI techniques that do not operate in context.
  5. 5.
    Finally, we conduct ablation studies of our approach, examining, for instance, alternative diffusion objectives and Gaussian approximations in place of flow matching, the model’s performance on out-of-distribution data, and the impact of problem dimensionality.
@@ 2 Related Work
Beyond the perspective of prior-data fitted networks, the contribution of this work can be summarized from the viewpoints of recent work on in-context learning, amortized Bayesian inference, and, simulation-based inference.
@@
ICL is a special case of meta-learning (Hospedales et al., 2021) characterized by using a large pre-trained model in order to learn from a context dataset without explicitly updating task-specific parameters. Several recent lines of work investigate the in-context learning capabilities of transformers (Garg et al., 2022; Ahuja et al., 2023; Wang et al., 2024; Chan et al., 2022).
Garg et al. (2022) show that a model similar to GPT-2 can implicitly implement various interesting function classes in context. More specifically, the model learns to reproduce the predictions of different statistical models such as (sparse) linear functions, decision trees, and even two-layer neural networks. This approach can be extended to multiple families of functions and even mixtures of tasks (Ahuja et al., 2023). Kirsch et al. (2022) investigate ICL as a general principle for meta-learning. However, the results by Garg et al. (2022) and Ahuja et al. (2023) are restricted to relatively small problem scales and scalar-valued predictions instead of multivariate posterior distributions. Additionally, the experiments are conducted exclusively on simulated data. In contrast, our results show that (large) transformer models can effectively learn multivariate posterior distributions over latent variables in context on real-world datasets. Furthermore, the focus on latent-variable models naturally steers our investigation toward unsupervised in-context learning, where the primary objective is to uncover the underlying structure of the data rather than to make predictions based on input-target paris presented in the context.
Concurrently, Mittal et al. (2025a) conduct a comparative analysis of amortized in-context Bayesian posterior estimation methods, ablating over different optimization objectives and architectural choices, and also reporting results on out-of-distribution performance and flow-matching methods. They focus on evaluating the posterior mean and downstream predictive performance, whereas we evaluate full posterior distributions. In an analogous setup, Mittal et al. (2025b) assess the effectiveness of learning point estimates versus learning entire distributions in context for the goal of predictive performance.
@@
Amortized inference is a central paradigm in the field of variational inference (Kingma, 2013; Zhai et al., 2018; Kim et al., 2018; Margossian & Blei, 2023). A commonly used idea here is to model the posterior distribution $P^{\bm{z}|\bm{x}}$ of latent variables $\bm{z}$ given a dataset $\bm{x}$ via a factorized density $p(\bm{z}|\bm{x})\approx\prod_{j=1}^{K}q_{\theta}(\bm{z}_{j}|h_{\bm{\phi}}(\bm{x}_{j}))$. In contrast to our more general assumption, each datapoint $\bm{x}_{j}$ in $\bm{x}$ is assumed to have a corresponding latent variable $\bm{z}_{j}$. While the parameter $\theta$ determines global aspects of the variational distribution, the function $h_{\bm{\phi}}$ is shared for all $\bm{x}_{j}$ and thus amortized across data $\bm{x}$. Variational autoencoders (Kingma, 2013; Rezende et al., 2014) and neural processes (Garnelo et al., 2018a, b; Rudner et al., 2018) are important model classes based on amortized inference.
In comparison, our ICL approach amortizes its parameters on the level of datasets, such that a single functional relationship is learned for a set $\mathcal{D}\subset(\mathcal{X}\times\mathcal{Z})^{N}$ of datasets. From this point of view, $\mathcal{D}=\left\{\left(\bm{x}_{i},\bm{z}_{i}\right)\right\}_{i=1}^{N}$ comprising $N$ datasets $\bm{x}_{i}\in\mathcal{X}$ and the corresponding latent variables $\bm{z}_{i}\in\mathcal{Z}$ can be seen as a “meta-dataset” for which we perform amortized inference. This is similar in nature to the setup by Le et al. (2017), who use recurrent neural networks to “compile” inference based on execution traces of probabilistic programs by training on simulated data.
Unlike in amortized variational inference, we do not use the notion of an evidence lower bound (Blei et al., 2017) or even the Kullback-Leibler divergence to learn the posterior, but utilize ideas that also appear in the context of simulation-based inference.
@@
Analogously to latent variable models, some scientific simulations, for instance in neuroscience or astrophysics (Fan & Markram, 2019; Schmit & Pritchard, 2018), allow to draw samples from the joint distribution $P^{\bm{x},\bm{z}}$ of data and latent variable of interest. Amortized posterior inference in this context is referred to as simulation-based inference (SBI; Cranmer et al., 2020). Several recent approaches focus on using neural networks to directly infer aspects of the likelihood $p(\bm{x}|\bm{z})$, the posterior $P^{\bm{z}|\bm{x}}$ or the joint distribution $P^{\bm{x},\bm{z}}$. More specifically, techniques based on discrete normalizing flows (Dax et al., 2021) or flow-matching (Wildberger et al., 2024) are used to approximate the posterior $P^{\bm{z}|\bm{x}}$, while (Gloeckler et al., 2024) propose to use a transformer-based diffusion model in order to approximate the joint distribution $P^{\bm{x},\bm{z}}$. In recent work, Vetter et al. (2025) directly a pre-trained TabPFN to auto-regressively sample $P^{\bm{z}|\bm{x}}$ leveraging ICL.
From a simulation-based inference viewpoint, we demonstrate that sample-based posterior estimation can be used for full Bayesian inference in complex scenarios arising in commonly used latent variable models, and demonstrate the effectiveness of this approach on real-world datasets.
@@ 3 In-Context Learning for Full Bayesian Inference
We propose a simple and effective solution based on ideas from ICL, which can be seen as conducting amortized inference on a dataset level. Training a model on a potentially unlimited amount of synthetic datasets yields an in-context learner that can not only approximate a vast, almost arbitrarily large, class of distributions but is also highly efficient when used for sampling. Furthermore, this does not incur the same issues with overly or insufficiently flexible distribution assumptions that are present in VI. More specifically, empirical results show that a major strength of TabPFN, for instance, is its ability to adapt flexibly to the complexity of the problem at hand, thus removing the need for extensive hyperparameter tuning (Hollmann et al., 2022).
In the following, we describe a sufficient general condition, as well as a specific framework that allows to train probabilistic in-context learners on simulated data.
@@
The idea underlying the proposed approach is founded on two observations relating to full Bayesian inference and the working principle of PFNs: First, many Bayesian models have a generative formulation that allows the simulation of arbitrarily large amounts of training samples from the joint distribution $P^{\bm{x},\bm{z}}$. We assume that samples from $P^{\bm{x},\bm{z}}$ comprise a dataset $\smash{\bm{x}=\left\{\bm{x}_{j}\right\}_{j=1}^{K}}$ containing $K$ samples $\bm{x}_{j}\in\mathcal{X}$ and a corresponding (latent) variable $\bm{z}\in\mathcal{Z}$. [^2] We do not assume any specific form of $\bm{z}$. That is, there can be a single $\bm{z}_{j}$ associated with each data point $\bm{x}_{j}$ in $\bm{x}$, but the case where a single “global” $\bm{z}$ governs the behavior of each $\bm{x}_{j}$ in $\bm{x}$ is equally included in this notation. This joint distribution $P^{\bm{x},\bm{z}}$ corresponds to the “prior” in PFNs and allows the training of a large neural network that implicitly learns to perform Bayesian inference. Second, Bayesian inference is especially useful for smaller datasets $\bm{x}$ that can be processed in a single forward pass. This makes an entire dataset a viable context for Bayesian ICL.
@@
More specifically, the central goal is to develop a method allowing to infer the posterior distribution $P^{\bm{z}|\bm{x}}$ of latent variables $\bm{z}\in\mathcal{Z}$, given observations $\bm{x}\in\mathcal{X}$ using ICL. From a supervised-learning perspective, we thus aim to directly learn the mapping $f_{0}:\mathcal{X}\rightarrow\mathcal{M}(\mathcal{Z}),\bm{x}\mapsto P^{\bm{z}|\bm{x}}$, where $\mathcal{M}(\mathcal{Z})$ is the space of all probability measures. Therefore, we want a model $f_{\theta}(\bm{x})=Q^{\bm{z}|\bm{x}}_{\theta}$ for the posterior to be as close as possible to the true posterior $P^{\bm{z}|\bm{x}}=f_{0}(\bm{x})$. We measure “closeness” w.r.t. some divergence $d:\mathcal{M}(\mathcal{Z})\times\mathcal{M}(\mathcal{Z})\to[0,\infty)$. When considering the expected divergence over data samples $\bm{x}\sim P^{\bm{x}}$, this gives rise to the following objective: $\mathcal{R}_{\theta}\vcentcolon=\mathbb{E}_{\bm{x}\sim p(\bm{x})}\left[d\left(f_{\theta}(\bm{x}),f_{0}(\bm{x})\right)\right]$, which can also be directly expressed as
$\mathcal{R}_{\theta}=\mathbb{E}_{\bm{x}\sim p(\bm{x})}\left[d\left(Q^{\bm{z}|\bm{x}}_{\theta},P^{\bm{z}|\bm{x}}\right)\right].$ (1)
@@
Note that we use the notion of a divergence $d$ loosely to refer to any measure of similarity of two distributions. Although $\mathcal{R}_{\theta}$ itself is usually intractable, specific choices of $d$ and the use of the joint distribution $P^{\bm{x},\bm{z}}$ make Equation 1 accessible via
@@
Let $d(Q_{\theta}^{\bm{z}|\bm{x}},P^{\bm{z}|\bm{x}})=\displaystyle\int\gamma\left(Q_{\theta}^{\bm{z}|\bm{x}}\right)dP^{\bm{z}|\bm{x}}$ for some measurable functional $\gamma:\mathcal{M}(\mathcal{Z})\rightarrow\mathbb{R}$.